The perimeter of a rectangle is 48 inches. The width of the rectangle is 2 more than half the length
Answer by addingup(3677)

Perimeter of this rectangle is:
2W+2L= 48 and the problem says that W=2+1/2L, so we substitute:
2L+2(2+1/2L)= 48
2L+4+L= 48 Now add L's on left and subtract 4 on both sides:
3L= 44 Divide both sides by 3:
L= 14.66
Check: 2(14.66)+ 2(2+7.33)= 48; 29.32+18.66= 48 Correct.